OBSERVATION OF H_{3}^{+} IN THE DIFFUSE INTERSTELLAR MEDIUM

\underline{BENJAMIN J. McCALL^{a}}, Department of Chemistry, The University of Chicago, Chicago, IL 60637; THOMAS R. GEBALLE, Joint Astronomy Centre, Hilo, HI 96720; KENNETH H. HINKLE, National Optical Astronomy Observatories, Tucson, AZ 85726; TAKESHI OKA, Department of Astronomy and Astrophysics, The University of Chicago, Chicago, IL 60637.

H_{3}^{+}, the cornerstone molecule of interstellar chemistry, was first detected in interstellar space in 1996^{b} in the dense molecular clouds GL2136 and W33A. We have now extended this ion’s diagnostic powers to the diffuse interstellar medium with the detection of H_{3}^{+} along the line of sight to the visible star Cygnus OB2 No. 12^{c}. Three rovibrational transitions in the 3.7 μm region were observed using the CGS4 infrared spectrometer at the United Kingdom Infrared Telescope and the high-resolution Phoenix spectrometer at the Kitt Peak National Observatory.

We have developed a simple chemical model of interstellar chemistry which describes the abundance of H_{3}^{+} in both diffuse and dense clouds. The application of this model to Cygnus OB2 No. 12 shows that this line of sight has H_{3}^{+} number density [H_{3}^{+}] ∼ 4 × 10^{−7} cm^{−3}, effective path length L ∼ 300 pc, and hydrogen number density [H] + 2[H_{2}] ∼ 20 cm^{−3}. The H_{3}^{+} absorptions also provide an estimate of the effective kinetic temperature of the medium T ∼ 27 K.

aSupported by the Fannie and John Hertz Foundation

bT.R. Geballe and T. Oka, Nature \underline{384}, 334 (1996)

cB.J. McCall, T.R. Geballe, K.H. Hinkle, and T. Oka, Science, in press